OpenStudy (anonymous):

A triangular prism has a surface area of 192 ft^2. The area of the base is 6 square feet,and the perimeter of the base is 12 feet. What is the height of the prism? a. 15.5 ft b. 28 ft c. 15 ft d. 17 ft. Explain. Thanks.

OpenStudy (anonymous):

192=(6^2)h of the prism

OpenStudy (anonymous):

Um, how did you get that?

OpenStudy (anonymous):

i believe, i dont see what perimenter has to do with it

OpenStudy (anonymous):

well the base is one side times the other if it is a square base

OpenStudy (anonymous):

wait srry i thought is was area not surface hang on

OpenStudy (anonymous):

Hmmm.

OpenStudy (anonymous):

Let me calculate the height then based on what you told me.

OpenStudy (anonymous):

I got h to be 5.33333333

OpenStudy (anonymous):

ok hang onTriangular based prism Base shape: Triangle: base 'b', height 'h', and sides S1, S2 and S3 Area of base: ½b×h Perimeter of base: S1+ S2 + S3 So, the formula is Surface area= 2*(area of base) + (base perimeter)*(prism height) Or Surface area = bh + (S1+ S2 + S3)H where b=Triangle base h=Triangle height S1=Side 1 of triangle S2=Side 2 of triangle S3=Side 3 of triangle H=Prism height

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@mzcandy216

OpenStudy (anonymous):

@niduj1234

OpenStudy (anonymous):

So, the formula is Surface area= 2*(area of base) + (base perimeter)*(prism height) Or Surface area = bh + (S1+ S2 + S3)H where b=Triangle base h=Triangle height S1=Side 1 of triangle S2=Side 2 of triangle S3=Side 3 of triangle H=Prism height the answer is B

OpenStudy (anonymous):

You used the theorem?

OpenStudy (anonymous):

yeah what did u want me 2 use

OpenStudy (anonymous):

How did you use the theorem?

OpenStudy (anonymous):

her formula is most accurate

OpenStudy (anonymous):

I would like to know how you used the theorem to find the answer.

OpenStudy (anonymous):

what theorem

OpenStudy (anonymous):

you used the pythagorean theorrem yes?

OpenStudy (anonymous):

no, that is with right triangles and neccesarily with this equation, why do u want to use it?

OpenStudy (anonymous):

is this what ur looking 4

OpenStudy (anonymous):

Triangular Prism Definition: A Triangular Prism is a polyhedron made of a triangular base, a translated copy, and 3 faces joining corresponding sides. Triangular Prism Formula : Area of Base(A) = ½ * a * b Perimeter of Base(P) = s1 + s2 + s3 Surface Area of Prism = ab + (s1 + s2 + s3)h = ab + Ph Volume of Prism = ½ * a * b * h = Ah

OpenStudy (anonymous):

I was just wondering how she/he got B.

OpenStudy (anonymous):

Can you do a drawing working the problem out? I need to show work on the paper.

OpenStudy (anonymous):

(Use the draw button)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

can i hvae a medal mz?

OpenStudy (anonymous):

Ummm..... I cannot give two medals on the same question.

OpenStudy (anonymous):

well this what im going to im going to use an different equation were my numbers at u put ur numbers there

OpenStudy (anonymous):

I cannot understand you.

OpenStudy (anonymous):

im going to do another problem and u just put ur numbers in same format just different numbers

OpenStudy (anonymous):

Ummm...Okay.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@ivanmlerner

OpenStudy (anonymous):

The base of a right triangular prism is a right triangle whose legs are 3 inches and 4 inches. If the height of the prism is 8 inches, find the Lateral Surface Area (LA) of the prism.We write the Pythagorean Theorem\[we use the pythgortherm c^2 =a^2+b^2,,,,,SUBMITE the legs of \triangle C^2=3(2) ,,,,,, the \square c^2=9+16 simpilfyc^2=25 the \square \root is 5 use sumin like this \to get threw ur promblem like i saud sub your problem \in and see if u get b like i d\]

OpenStudy (anonymous):

I do not understand...

OpenStudy (anonymous):

How do you find the height?

OpenStudy (anonymous):

Base shape: Triangle: base 'b', height 'h', and sides S1, S2 and S3 Area of base: ½b×h Perimeter of base: S1+ S2 + S3 So, the formula is Surface area= 2*(area of base) + (base perimeter)*(prism height) Or Surface area = bh + (S1+ S2 + S3)H where b=Triangle base h=Triangle height S1=Side 1 of triangle S2=Side 2 of triangle S3=Side 3 of triangle H=Prism height

OpenStudy (anonymous):

Surface area= 2*(area of base) + (base perimeter)*(prism height

OpenStudy (anonymous):

Thanks,Mzcandy. I got the answer to be C.

OpenStudy (anonymous):

Taking the area of the bases from the total area get get 180 square foot. Now, since the height is the same for all rectangles of the side, we get that the total area without the bases is:\[ah+bh+ch=180\]where a b and c are the sides of the base.\[(a+b+c)h=180\]But a+b+c is the perimeter given so:\[12h=180\]\[h=15\]

OpenStudy (anonymous):

I believe this way its easier

OpenStudy (anonymous):

Hmmm. Well, thqnks:)